What is the present value of an annuity if the size of each payment is Birr 100 payable at the 

end of each quarter for 2 years and the interest rate is 8% compounded quarterly?


1
Expert's answer
2022-01-16T14:50:09-0500

Solution:

Given, P=100, t=2, n=4 (as quarters), r=8% compounded quarterly.

We know that, "PV=P\times\dfrac{1-(1+\frac rn)^{-nt}}{\frac rn}"

Putting given values,

"PV=100\times\dfrac{1-(1+\frac {0.08}4)^{-8}}{\frac {0.08}4}
\\=100\times\dfrac{0.1465096}{0.02}
\\=732.548\ birr"